eliminate(thita) from the equations
atan(thita)+bcot2(thita)=c
and acot(thita)-btan2(thita)=c

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
13 years ago

Dear ravi

 

atanθ+bcot2θ=c  ........1

acotθ-btan2θ=c  ..........2


subtract equation

a(cotθ -tanθ)=b(cot2θ+tan2θ)

a cos2θ/cosθsinθ = b/cos2θ sin2θ

2a/tan2θ  = 2b/sin4θ

sin4θ /tan2θ =b/a

2tan2θ/(1+tan22θ).tan2θ =b/a

let T =tan2θ

2T/(1+T2)T =b/a

T2 =(2a-b)/b  ....................3

and from the equation 1 and 2

c-b/T =atanθ

c+bT =acotθ

multiply

(c-b/T)(c+bT) =a2

c2 -b2 +bc(T-1/T) =a2

c2 -b2 +bc(T2-1)/T =a2

put the value of T

c2 -b2 +bc(2a-2b)/b  .√b/(√2a-b) =a2

2c(a-b)√b =(a2+b2-c2)√(2a-b)

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.